### v1

# “My life is miserable, have to sign 500 autographs everyday”: Exposing Humblebragging, the Brags in Disguise

Abstract
Humblebragging is a phenomenon where individuals present self-promotional statements under the guise of modesty or complaints. For example, a statement like, “Ugh, I can’t believe I got promoted to lead the entire team. So stressful!”, subtly highlights an achievement while pretending to be complaining. Detecting humblebragging is important for machines to better understand the nuances of human language, especially in tasks like sentiment analysis and intent recognition. However, this topic has not yet been studied in computational linguistics. For the first time, we introduce the task of automatically detecting humblebragging in text. We formalize the task by proposing a 4-tuple definition of humblebragging and evaluate machine learning, deep learning, and large language models (LLMs) on this task, comparing their performance with humans. We also create and release a dataset called HB24, containing 3,340 humblebrags generated using GPT-4o. Our experiments show that detecting humblebragging is non-trivial, even for humans. Our best model achieves an F1-score of 0.88. This work lays the foundation for further exploration of this nuanced linguistic phenomenon and its integration into broader natural language understanding systems.
“My life is miserable, have to sign 500 autographs everyday”: Exposing Humblebragging, the Brags in Disguise
Sharath Naganna1, Saprativa Bhattacharjee1, Pushpak Bhattacharyya1, Biplab Banerjee2, 1Computer Science and Engineering, Indian Institute of Technology Bombay, India 2Centre of Studies in Resources Engineering, Indian Institute of Technology Bombay, India {sharathhn, saprativa, pb}@cse.iitb.ac.in, bbanerjee@iitb.ac.in
$\ast$\$$\ast$\$footnotetext: Equal Contributions
1 Introduction
Humblebragging is a nuanced socio-linguistic phenomenon where individuals subtly boast about their achievements, possessions, or qualities while disguising their self-promotion with expressions of complaint or humility. The term was coined by American comedian Harris Wittels in 2010, who later published a book on the phenomenon titled Humblebrag: The Art of False Modesty Wittels (2012).
To address why people resort to humblebragging, Trivedi and Srinivas (2019) analyzed its rising popularity, particularly on social media, from the dual perspective of need theories (Maslow’s Hierarchy of Needs, Maslow and Lewis 1987; see footnote 2) and the Hubris Hypothesis Hoorens et al. (2012). Humblebragging allows a person to satisfy the dual needs of belonging (level 3) through humility and self-esteem (level 4) through bragging. Moreover, the hubris hypothesis states that addressees prefer implicit self-superiority claims over explicit ones. Straightforward bragging also violates the maxims of modesty Leech (1983) and self-denigration Gu (1990), as noted by Zuo (2023). Humblebragging thus serves as a subtle way to convey positive information about oneself, often through expressions of modesty or self-deprecation. Even though it is commonly observed on social media (Twitter, Reddit, and Instagram), many people also use it in real-life conversations Wittels (2012); Sezer et al. (2018); Trivedi and Srinivas (2019). Examples of humblebrags are provided in Table 1.
Figure 1: Why do people resort to humblebragging? The answer lies in the Maslow’s Hierarchy of Needs [^2] Image source: https://www.simplypsychology.org/maslow.html. Humblebragging satisfies needs of belonging and self-esteem simultaneously.
Humblebrags Mask type
I can’t believe they’d give an idiot like me a phd lol Modesty
Being in demand means disappointing 95% of people 95% of the time. I have yet to learn how to overcome this. Modesty
For the 3rd time in 3 years I’ve been asked to speak at Harvard, but I’ve yet to speak at my alma mater. What’s a girl gotta do @MarquetteU? Complaint
Will Twitter be available for me in Paris, milan, or the Maldives? I hope so bc it won’t in hong Kong or Singapore Complaint
Table 1: Examples of humblebrags. Each instance of a humblebrag consists of a brag masked by either complaint or modesty. The brags are in red while the masks are in blue.
Humblebragging is closely related to other forms of figurative language, such as sarcasm Gibbs (1986) and irony Garmendia (2018). All three rely on verbal incongruity, where there is a contrast between what is being said and what is actually meant. While extensive research exists on the computational modeling of sarcasm Bhattacharyya and Joshi (2017); Riloff et al. (2013); Cai et al. (2019) and irony Zeng and Li (2022); Van Hee et al. (2018); Barbieri and Saggion (2014), to the best of our knowledge, this work is the first to explore humblebragging in computational linguistics
Detecting humblebragging automatically is crucial for a nuanced understanding and processing of human language. This is necessary for enhancing accuracy of tasks such as sentiment analysis Zhang et al. (2024), intent recognition Lamanov et al. (2022), emotion recognition Li et al. (2022) and dialogue understanding Gao et al. (2024). In applications like social media monitoring and customer feedback analysis, it helps differentiate genuine complaints Singh et al. (2023) from brags disguised as complaints. Moreover, this capability is also valuable for researchers in the humanities and social sciences.
In this paper, we formally introduce the task of humblebragging detection and present a curated dataset that combines existing resources with synthetic data generated using GPT-4o. By addressing this previously unexplored area, our work bridges the gap between computational linguistics and other disciplines in the study of humblebragging.
Our contributions are:
  1. 1.
    Introducing the task of automatic humblebragging detection from text to the computational linguistics community by proposing a 4-tuple definition of humblebragging for streamlining its processing (section 3).
  2. 2.
    Benchmarking various machine learning, deep learning, and state-of-the-art large language model techniques on the task of automatic humblebragging detection from text (section 5 and section 6).
  3. 3.
    Releasing a new dataset named HB24 on humblebragging detection, containing 3,340 humblebrags, to enable further research on the task (section 4).
2 Related Work
In Psychology and Other Disciplines
Humblebragging has been extensively studied in psychology. Sezer et al. (2018) examined its effects on audiences, showing that humblebragging is ubiquitous in daily interactions, with 70% of humblebrags falling into the complaint-masked variety. Trivedi and Srinivas (2019) explained the widespread use of humblebragging through the dual perspectives of need theories and the hubris hypothesis, and also provided a contextual framework for understanding it. Other notable studies include Sezer et al. (2015); Vranka et al. (2017); Luo and Hancock (2020). Beyond psychology, humblebragging has been explored in disciplines like tourism research Chen et al. (2020); Yan et al. (2024), pragmatics Lin and Chen (2022); Zuo (2023); Han et al. (2024), and advertising Paramita and Septianto (2021).
Sarcasm and Irony
Both sarcasm and irony have been extensively studied in computational linguistics over the past two decades. Joshi et al. (2015) demonstrated how incongruences can enhance sarcasm detection, while Joshi et al. (2017a) categorized detection methods, benchmark datasets, and evaluation metrics. Recently, Gole et al. (2023) explored the use of large language models for sarcasm detection. Beyond detection, Joshi et al. (2017b) proposed a hybrid rule-based and statistical approach for identifying sarcasm targets, complemented by transformer-based methods like BERT, as shown by Parameswaran et al. (2021). For irony, Zeng and Li (2022) provided a comprehensive survey on computational approaches. Hernández-Farías et al. (2015) evaluated traditional machine learning models for irony detection using sentiment scores, while Wen et al. (2023) introduced the Retrieval–Detection Method for Verbal Irony (RDVI), leveraging open-domain resources for enriched detection.
Bragging and Humility
A closely related area of research is the detection and processing of bragging Alfano and Robinson (2014) and humility Snow (1995) as standalone tasks. Jin et al. (2022) introduced bragging classification to the computational linguistics community and released a public dataset, while Jin et al. (2024) conducted a large-scale study of bragging behavior on Twitter. For humility, Guo et al. (2024) explored LLM-based techniques for measuring humility in social media posts. Additionally, Danescu-Niculescu-Mizil et al. (2013); Firdaus et al. (2022); Srinivasan and Choi (2022) have examined politeness in computational contexts.
Synthetic Data Generation
The language generation capabilities of LLMs have created opportunities for generating synthetic data. Long et al. (2024) provide a comprehensive survey of synthetic data generation, curation, and evaluation, while Li et al. (2023) explore the potential and limitations of using LLMs for this purpose. Synthetic data generated by language models has been applied to various text classification tasks Chung et al. (2023); Sahu et al. (2022); Ye et al. (2022); Yoo et al. (2021).
Classification as Generation
Finally, there has been a growing trend towards performing classification tasks by posing them as generation tasks. This approach to text classification is particularly relevant for leveraging decoder-based large language models in classification settings, where text generation mechanisms complement traditional methods. For instance, LLMs have been employed as zero-shot Gretz et al. (2023) and few-shot Mirza et al. (2024) text classifiers. Moreover, Saunshi et al. (2021) provides mathematical insights into modeling classification tasks as text completion tasks
3 Task Formulation
In this section, we formally define humblebragging and outline the associated tasks.
3.1 Formal Definition of Humblebragging
Similar to the 6-tuple definition of sarcasm proposed by Ivanko and Pexman (2003), we define humblebragging as a 4-tuple to systematically capture its key components and underlying structure:
$HB=\langle B,BT,HM,MT\rangle$ (1)
where:
  • $B$: Brag – The segment of the text that explicitly conveys the act of bragging.
  • $BT$: Brag Theme – The overarching theme or specific category of the brag embedded within the statement. Categories are listed in Appendix G.
  • $HM$: Humble Mask – The segment of the text that adopts a modest or self-deprecating tone to obscure or mitigate the act of bragging.
  • $MT$: Mask Type – Specifies whether the humble mask adopts a modest tone or a self-deprecating approach.
For instance, in the following statement:
"Ugh, I can’t believe I got promoted to lead the entire team. So stressful!"
  • $B$: "I can’t believe I got promoted to lead the entire team.";
  • $BT$: Performance at work;
  • $HM$: "Ugh," and "So stressful!";
  • $MT$: Complaint.
3.2 Binary Classification for Humblebragging Detection
In the context of humblebragging detection, the task is to classify a given text $x$ as either $C_{\text{HB}}$ (Humblebragging) or $C_{\text{Non-HB}}$ (Non-Humblebragging). The process involves generating text encodings from the input, which are then used for classification. For further details, we refer the reader to Appendix A.
3.3 Classification as a Sentence Completion Task with Yes/No Questions
Though decoder models are primarily designed for language generation tasks, their ability to predict the next token in a sequence makes them adaptable to various natural language understanding tasks, including classification. Humblebragging classification can be reformulated as a Yes/No question-answering task, where the model determines whether the input text contains humblebragging. This approach leverages the natural language understanding capabilities of pre-trained language models to classify text.
Framework
The input text $x$ is transformed into a prompt structured as a question. For example:
$<question><x><answer>$
The model is given the $question$ and $x$ as the input prompt and is expected to generate a text completion for the $answer$. The $answer$ is then analyzed to determine whether it contains the required word.
  • If $answer$ contains "Yes", the input is classified as $y=1$ (Humblebragging).
  • If $answer$ contains "No", the input is classified as $y=0$ (Non-Humblebragging).
4 Dataset
As there were no existing datasets for the task of humblebragging detection, we propose HB24, a well-balanced collection of humblebrag and non-humblebrag texts, comprising both human-written and synthetic samples. Due to the limited availability of quality humblebrags, we leverage the capabilities of large language models to generate human-like examples, augmenting the existing data and enhancing the corpus for training classification models. The synthetic data is used for training, while the trained model’s performance is evaluated on human-written samples. Table 2 presents the dataset statistics.
Humblebrag Non-humblebrag
Train Samples 3340 5431
Min 6 1
Max 47 68
Avg 15.98 16.41
Test Samples 558 576
Min 1 6
Max 70 47
Avg 19.55 17.5
Table 2: Dataset statistics. Min, Max and Avg refers to the minimum, maximum and average number of words respectively.
4.1 Human-Written Humblebrags
Wittels (2012) presents a curated collection of high-quality humblebrag tweets, categorized into themes such as wealth, first-class travel, workplace achievements, celebrity status, and more. These tweets form the positive class within our test set.
4.2 Synthetic Humblebrags
The humblebrags in the training set consist entirely of synthetic tweets generated using GPT-4o through zero-shot and few-shot prompting. The prompt template follows a format similar to that of Li et al. (2023). In the following sections, we discuss the prompts used for generating synthetic humblebrags.
4.2.1 Zero-Shot Prompts
In the zero-shot generation setup, we used two types of prompts. In the General Prompt, we did not explicitly define humblebragging; instead, we asked the model to generate tweets that subtly mention various achievements. In the Prompt with Themes, we provided a formal definition of humblebragging along with the themes (Appendix G) outlined in Wittels (2012). The prompts are provided in Appendix B.
4.2.2 Few-Shot Prompts
In the few-shot prompt setup, we modified the Prompt with Themes to include a few examples from each theme. We experimented with varying numbers of examples, starting with one and going up to five, and observed that increasing the number of examples did not improve the generation quality. Consequently, we settled on three examples per prompt for generating samples with few-shot prompts.
Prompt Type #Samples
General prompt 1100
Prompt with Themes 1304
Few-Shot with Themes 936
Total 3340
Table 3: Prompt type and the number of samples.
4.2.3 Post-Processing and Data Curation
After executing all three prompts, we generated a total of 11,000 synthetic samples containing humblebrags. Each sample was manually reviewed to assess its quality and relevance. From this pool, we filtered tweets from each prompt type, ensuring a balanced selection (see Table 3).
Train Test
Sarcasm 16% 12%
Humblebrag 38% 49%
Irony 15% 11%
Complaints 14% 10%
Neutral 14% 15%
Bragging 3% 4%
Table 4: Dataset composition.
4.3 Non-Humblebrags
Humblebrags are often confused with sarcasm and irony, as all involve an incongruence between the utterance and its intended meaning. To help the model distinguish these phenomena, we included sarcasm and irony as negative samples, alongside direct brags and straightforward complaints. Direct brags convey explicit self-promotion, while complaints reflect surface emotions often present in humblebrags.
Sarcastic samples were sourced from the SARC dataset Khodak et al. (2018), and ironic ones from SemEval-2018 Van Hee et al. (2018). Brags and complaints were taken from Jin et al. (2022) and Preoţiuc-Pietro et al. (2019), respectively. Neutral sentences, essential for improving class distinction Koppel and Schler (2006), were drawn from SemEval-2017’s sentiment analysis task Rosenthal et al. (2017). The dataset composition is shown in Table 4.
5 Experimental Setup
5.1 Machine Learning Classifiers
We conducted a grid search on various hyperparameters to identify the best combination for each classifier. For the Logistic Regression model, we set the maximum number of iterations (max_iter) to 100 to ensure convergence, the regularization strength (C) to 0.1 to control overfitting, and a fixed random state (random_state=42) for reproducibility. For the Support Vector Classifier (SVC), we used a radial basis function (RBF) kernel (kernel=‘rbf’) to capture non-linear relationships in the data and set the regularization parameter (C) to 10.
5.2 Encoder Models
Two transformer-based Vaswani et al. (2017) encoder models, BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019), were evaluated on the task of humblebragging classification. The experiments utilized the Adam optimizer, and 5-fold cross-validation was employed for hyperparameter tuning. Details of the hyperparameters and training configurations are provided in Appendix E.
5.3 Decoder Models
For the task of detecting humblebragging, decoder models were evaluated in three settings: zero-shot classification (Z), zero-shot with the 4-tuple definition of humblebragging (Z+D), and fine-tuning on the humblebragging dataset (F). In Z, models classify statements as humblebrag or not, relying only on the input. In Z+D, models were contextually grounded using the 4-tuple definition introduced in subsection 3.1. The prompts are mentioned in Appendix D.
We used open-source models namely, Qwen2.5-7B-Instruct Yang et al. (2024), Llama 3.1-8B-Instruct Dubey et al. (2024), Gemma 1.1 7B (IT) Team et al. (2024), Mistral-7B-Instruct-v0.3 Zheng et al. (2023a), and Vicuna-7B-v1.5 Zheng et al. (2023b) from Hugging Face’s pre-trained model repository [^3] https://huggingface.co/, along with proprietary models GPT-3.5 and GPT-4o. Model outputs were limited to two tokens, and each model was evaluated five times to record average metrics. GPT-3.5 and GPT-4o predictions were obtained via OpenAI’s API [^4] https://platform.openai.com/docs/overview. Hyperparameter details are provided in Appendix E.
5.4 Human Performance
To evaluate human performance, the test dataset was labeled by three independent annotators, including two with Masters degrees in Linguistics and Arts, and a final-year Masters student in Computer Applications. All were proficient in English and experienced in professional annotation. An initial meeting with the authors covered annotation guidelines and the scope of the task. The annotators first labeled a small subset of the dataset, and disagreements were discussed to ensure consistency. They then labeled the entire test set as yes or no, indicating the presence or absence of humblebragging. The annotators were compensated according to University norms.
5.5 Evaluation Metrics
We evaluate model performance using Macro Precision, Macro Recall, Macro F1-Score, and Accuracy, ensuring a comprehensive analysis of the results.
6 Results and Discussion
We present the results of all our experiments in Table 5. The table begins with the majority class baseline, which, in our case, involves predicting every sample as non-humblebrag.
6.1 Quantitative Analysis
Model Precision Recall F1-Score Accuracy
Baseline 0.25 0.50 0.34 0.51
Human 1 0.89 0.81 0.85 0.86
Human 2 0.86 0.81 0.84 0.84
Human 3 0.82 0.51 0.63 0.70
Average 0.86 0.71 0.77 0.80
Logistic Regression 0.68 0.58 0.53 0.59
SVM 0.72 0.61 0.56 0.62
BERT-Large-Uncased (F) 0.76 0.50 0.61 0.68
RoBERTa-Large (F) 0.91 0.62 0.74 0.78
GPT-4o (Z) 0.78 0.94 0.85 0.84
GPT-4o (Z+D) 0.91 0.85 0.88 0.89
GPT-3.5 (Z) 0.65 0.60 0.57 0.61
GPT-3.5 (Z+D) 0.76 0.75 0.75 0.75
Qwen2.5-7B-Instruct (Z) 0.82 0.35 0.49 0.64
Qwen2.5-7B-Instruct (Z+D) 0.85 0.50 0.63 0.71
Qwen2.5-7B-Instruct (F) 0.85 0.40 0.54 0.67
Mistral-7B-Instruct-v0.3 (Z) 0.55 0.96 0.70 0.60
Mistral-7B-Instruct-v0.3 (Z+D) 0.55 0.96 0.70 0.60
Llama-3.1-8B-Instruct (Z) 0.49 0.99 0.66 0.49
Llama-3.1-8B-Instruct (Z+D) 0.62 0.88 0.72 0.68
Llama-3.1-8B-Instruct (F) 0.87 0.72 0.79 0.81
Gemma-1.1-7b-it (Z) 0.57 0.57 0.57 0.57
Gemma-1.1-7b-it (Z+D) 0.53 0.83 0.65 0.56
Gemma-1.1-7b-it (F) 0.71 0.44 0.60 0.71
Vicuna-7b-v1.5 (Z) 0.60 0.28 0.38 0.55
Vicuna-7b-v1.5 (Z+D) 0.62 0.51 0.56 0.61
Table 5: Results of humblebragging classification. Z: zero-shot, Z+D: zero-shot with 4-tuple definition, F: fine-tuned. The best values are in bold.
Overall, the best-performing model in terms of F1-score is GPT-4o (0.88 F1), outperforming even the best human annotator (0.85 F1). We speculate this might be due to the vast linguistic and world knowledge acquired by such large parameter models. Among the three human annotators, one performed significantly worse than the others, indicating that the task is non-trivial and may even be difficult for some individuals. Moreover, two notable observations emerge from the results.
First, across all decoder models, Z+D versions consistently outperform their corresponding Z versions. This demonstrates that our 4-tuple definition aids the models in identifying instances of humblebragging. To further verify that the 4-tuple definition enhances the models’ understanding of humblebragging, we conducted additional experiments: (a) replacing the definition in the system prompt with a completely random gibberish text (Appendix D) and (b) replacing our 4-tuple definition with an alternative definition [^5] We refer to this as the textbook definition (TD). provided by Wittels in his book Wittels (2012).
A specific type of brag that masks the boasting part of a statement in a faux-humble guise. The false humility allows the offender to boast about their “achievements” without any sense of shame or guilt. Humblebrags are usually self-deprecating in nature.
The results presented in Figure 2 clearly demonstrate that our 4-tuple definition significantly improves the models’ performance as humblebrag classifiers.
Figure 2: Comparison of F1-scores across scenarios for decoder models: G (Gibberish), Z (Zero-Shot), Z+D (Zero-Shot with our proposed 4-tuple definition of humblebragging), and Z+TD (Zero-Shot with the textbook definition of humblebragging). The Z+D setting achieves the highest F1-scores across all models, demonstrating the effectiveness of the proposed 4-tuple definition in capturing humblebragging nuances.
Second, fine-tuning with our HB24 dataset improved the F1-scores of the majority of the models. Both encoder models and three decoder models (Llama, Gemma, and Qwen) showed significant gains from fine-tuning. Interestingly, fine-tuned RoBERTa outperformed all 7–8 billion parameter decoder models except for Llama (F). This highlights the superior classification capabilities of encoder-only models when a high-quality dataset for fine-tuning is available.
We illustrate in Figure 3 Llama’s progression from an F1-score of 0.66 in the zero-shot setting (Z) to an F1-score of 0.79 after fine-tuning (F), through confusion matrices. Llama (Z) primarily predicted the yes label for almost all samples. With our 4-tuple definition in Llama (Z+D), the model began to identify non-humblebrags, bringing more balance to the confusion matrix. After fine-tuning, Llama (F) became more proficient in identifying non-humblebrags while sacrificing some true positives.
Lastly, we observed anomalous behavior for Mistral and Vicuna, where fine-tuning resulted in a decrease in F1-scores. For Mistral, one possible explanation could be that its zero-shot (Z) performance in humblebragging classification was already much higher than all other models in its category and even exceeded that of the much larger GPT-3.5. In the case of Vicuna, the fine-tuned model generated random texts and emojis. Also we observed very high sensitivity to minimal changes in prompts in Z+D setting. The same prompt used for other models resulted in a drop in F1-score. We had to remove the last sentence from the system prompt (Appendix D) which improved the performance. This behavior of Vicuna demands further investigation.
(a) Llama (Z)
(b) Llama (Z+D)
(c) Llama (F)
Figure 3: Confusion matrices for Llama Z vs Z+D vs F. Gradual improvement in the model can be observed from left to right.
6.2 Qualitative Analysis
In this section, we analyze cases of agreement and disagreement between human annotators and models.
All humans and models correctly classified the following as humblebrags, as the brags were easy to identify with a clear distinction between the brag and the mask segments:
T1: In the limo riding to airport. Sucks being alone though
T2: just tried to pre-order my book. couldnt figure it out. did anyone try?
For the following humblebrags, humans classified them as no, while models classified them as yes:
T3: I forget. What airport do u fly into to get to Maui?
T4: I just had my first screaming girl encounter. She probably had me confused withsomeone else.
In the case of T3, the disagreement could stem from a lack of knowledge regarding Maui as an exotic travel destination. On the other hand, the phrase “screaming girl” in T4 might not have been understood by the annotators due to cultural differences. Some cultures might interpret the phrase at its surface level without delving into its deeper meaning.
For the following humblebrag, humans said yes, while models said no:
T5: The CNN-LA green room is a cold and lonely place at 7 on a Sunday morning.Funnily enough, CNN LA green room a cold and lonely place at 10 on a Monday too.
We hypothesize that the model might have been confused by the incongruity between “cold and lonely” and “Funnily enough,” interpreting it as sarcasm instead of a humblebrag.
For the following non-humblebrag, both humans and models classified it as yes:
T6: i decided to become my own boss to have more free time.. now i have no time left whatsoever.
T6 is a rare case where our assumption that humblebrags are not present in the datasets used to create our negative samples was violated. After encountering this example, we reviewed our dataset again to ensure no other such cases exist.
For the following non-humblebrag, models said no, but humans said yes:
T7: After years of fumbling around , I have finally found a skin care product that works for me . Well , at least for now ?
T8: TheoCorleone david_maclellan Shit! I better shut my stupid girly mouth because im so concerned about what men might think of me.
T7 and T8 represent classic cases of annotator bias, where annotators attempt to imagine a non-existent context and incorrectly classify the samples as belonging to the positive class. This bias arises because annotators, subconsciously influenced by their task, oversearch for humblebrags in the data they are tasked to annotate for humblebragging detection.
7 Conclusion and Future Work
We introduced the task of automatic humblebragging detection, formalized through our proposed 4-tuple definition. We benchmarked various machine learning, deep learning, and large language models on this task, providing a comparative analysis against human performance. We also demonstrate that our 4-tuple definition significantly improves the zero-shot capabilities of all decoder models. Additionally, we released a synthetic dataset, HB24, generated using GPT-4o, to facilitate further research. Our experiments and analysis reveal that detecting humblebragging is a challenging task, even for humans. This study lays the groundwork for exploring this intricate linguistic phenomenon and its integration into natural language understanding systems.
Future research could focus on developing improved models for classifying humblebragging, leading to a deeper understanding of this unique communication style. Another promising direction would be the generation of humblebrag captions for images. Lastly, machines could be trained to rewrite direct brags as humblebrags.
Limitations
The inherent subjectivity of humblebragging complicates the creation of universally agreed-upon labels, as even humans often struggle to classify such statements consistently. Additionally, while machine-generated texts are sophisticated and well-structured, they often lack the spontaneity and imperfections typical of human-authored texts. For instance, the model’s inability to use certain casual or curse words, as well as elongated words like sooooo or goood, which are often present in human-written humblebrags. This creates a mismatch when using synthetic datasets like HB24, which, despite being a valuable resource, may fail to fully capture the linguistic diversity and subtleties of real-world humblebragging, thereby limiting the generalizability of trained models. Moreover, the task itself remains underexplored, with no prior benchmarks or resources, making it difficult to contextualize results within the larger field of natural language understanding.
Appendix A Binary Classification
Text Encoding Generation
The input text $x$ is converted into a numerical representation $\mathbf{e}$ using any of the available encoding techniques. Formally:
$\mathbf{e}=f_{\text{encoder}}(x)$ (2)
where:
  • $x$: Input text.
  • $\mathbf{e}$: Encoded representation of the text, typically a fixed-dimensional vector.
  • $f_{\text{encoder}}$: The encoding function, such as TF-IDF, BERT or a similar pre-trained transformer.
This encoded representation $\mathbf{e}$ captures semantic and contextual information from the input text, enabling effective classification.
Binary Classification
Using the encoded representation $\mathbf{e}$, the model predicts the probability $\hat{y}\in[0,1]$ for the text belonging to the class $C_{\text{HB}}$. The true label $y$ is $y=1$ for humblebragging and $y=0$ otherwise.
The Binary Cross-Entropy (BCE) loss for this task is defined as:
$\mathcal{L}_{\text{BCE}}(y,\hat{y})=-\frac{1}{N}\sum_{i=1}^{N}\Big{(}y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\Big{)}$
where:
  • $N$: Total number of samples in the dataset.
  • $y_{i}$: True label for the $i$-th sample ($y_{i}\in\{0,1\}$), where $y_{i}=1$ indicates a humblebrag.
  • $\hat{y}_{i}$: Predicted probability that the $i$-th sample is a humblebrag ($\hat{y}_{i}\in[0,1]$).
Objective
The model is trained to minimize $\mathcal{L}_{\text{BCE}}$ over the dataset, improving its ability to accurately classify texts as humblebragging or non-humblebragging.
Appendix B Data Generation Prompts
General Prompt
You are now a person about to humblebrag
about your recent achievement to attract
peoples attention and make them praise
you.But you cant state the obvious. You
have to present it in such a way that it
sounds like a complaint without reducing
the importance of the achievement.
There should be a strong incongruence.
Make sure these are tweets, and keep the
tone casual. Be specific about your
achievements and use diverse topics.
Do not use topics already generated,
and do not follow a pattern for
beginning the text.
Prompt with Themes
Here is the definition of humblebragging:
a specific type of brag that masks the
boasting part of a statement in a
faux-humble guise. The false humility
allows the offender to boast about their
"achievements" without any sense of shame
or guilt. Humblebrags are usually
self-deprecating in nature.
Now, you are a person who is about to
humblebrag on Twitter with the theme
<theme> and it should sound casual.
Use the above definition and generate
humblebrags.
Appendix C Pre-Processing
Unlike other types of textual data, tweets are often informal and unique in their composition. They frequently include emojis and emoticons, which add emotional or contextual cues. Additionally, tweets commonly feature elongated words (e.g., soooo or goood) and repeated characters for emphasis or emotional expression. While modern tokenizers utilized by pre-trained networks are designed to handle these emojis and elongated words effectively, traditional machine learning algorithms often struggle with such unconventional text patterns.
For machine learning classifiers, we incorporated existing pre-processing techniques in two distinct phases. In the first phase, we replaced all emojis with their corresponding verbal explanations, as suggested by Singh et al. (2019), to retain the semantic information. In the second phase, we utilized ekphrasis Baziotis et al. (2017), a specialized text pre-processing tool, to handle hashtags, elongated and repeated words, URLs, and numeric information. The output was further preprocessed by removing stop words and punctuation, constructing unigram and bigram tokens, limiting the vocabulary to the top 10,000 tokens by frequency, and requiring each token to appear in at least two training samples. The resulting numerical representations, created using TF-IDF, were used as inputs for machine learning classifiers such as Support Vector Machine and Logistic Regression.
Appendix D Inference Prompts
User Prompt
### Question: Is this a humble brag?
Answer in yes or no only.
### Statement: {data_point[’text’]}
### Answer:
System Prompt
A humble brag comprises the following
components:
1. Brag:
- The segment of the text that
explicitly conveys the act of
bragging.
2. Brag Theme:
- The overarching theme or specific
category of the brag embedded within
the statement.
- Possible categories include:
- Looks and Attractiveness
- Achievements
- Performance at Work
- Money and Wealth
- Intelligence
- Personality
- Social Life
- Miscellaneous
3. Humble Mask:
- The element of the text that adopts
a modest or self-deprecating tone to
obscure or mitigate the act of
bragging.
4. Mask Type:
- Specifies whether the humble mask
adopts a modest tone or a
self-deprecating approach.
Now you are about to classify if a given
sentence is a humble brag or not using
the above definition.
Random Gibberish Prompt
A Quantum Pancake involves the following
components:
1. Flapjack Fluctuation:
- Analyze the positive or negative
curvature of the syrupy timeline
as the pancake flips through
space-time.
2. Stack Dynamics:
- Identify the structural
integrity of the pancake layers
and their inter-dimensional
fluffiness coefficient.
3. Butter Singularities:
- Highlight the concentrated
points of creamy chaos where the
butter both exists and does not
exist simultaneously.
4. Maple Entanglement:
- Describe the sticky phenomena
where the syrup defies Newtonian
logic to connect pancakes across
parallel brunch universes.
Now you are about to classify if a given
sentence is a humble brag or not using
the above definition.
Appendix E Hyperparameters
The hyperparameters for the experiments are as follows.
For encoder models, BERT-Large-Uncased (340M) was trained with a learning rate of 5e-3, batch size of 16, and 4 epochs, while RoBERTa-Large (355M) used a learning rate of 5e-4, batch size of 32, and 5 epochs. Both models shared a maximum sequence length of 128, a weight decay of 0.01, a warmup ratio of 0.1, and gradient clipping at 1.0.
For decoder models, fine-tuning was conducted using LoRA (Low-Rank Adaptation) with a scaling factor (lora_alpha) of 8, rank (r) of 16, and targeted attention modules (q_proj and k_proj). The dataset was split into 80% training and 20% validation with a random seed of 42. Training employed a cosine learning rate scheduler with a learning rate of 1e-5, weight decay of 0.01, and 10 warmup steps, over 4 epochs with a batch size of 8 and gradient accumulation steps of 4. The maximum sequence length was set to 512 tokens, and the SFTTrainer was utilized for efficient fine-tuning.
All experiments were conducted on NVIDIA A100-SXM4-80GB GPUs, utilizing approximately 300 GPU hours in total.
Appendix F Confusion Matrices
(a) Qwen (Z)
(b) Qwen (Z+D)
(c) Qwen (F)
Figure 4: Confusion matrices for Qwen Z vs Z+D vs F.
(a) Gemma (Z)
(b) Gemma (Z+D)
(c) Gemma (F)
Figure 5: Confusion matrices for Gemma Z vs Z+D vs F.
(a) Mistral (Z)
(b) Mistral (Z+D)
Figure 6: Confusion matrices for Mistral Z vs Z+D.
(a) Vicuna (Z)
(b) Vicuna (Z+D)
Figure 7: Confusion matrices for Vicuna Z vs Z+D.
(a) BERT (F)
(b) RoBERTa (F)
Figure 8: Confusion matrices for BERT (F) and RoBERTa (F).
Appendix G Humblebrag Categories
Category Sezer et al. (2018) Theme Wittels (2012)
Looks and Attractiveness Ugh, Being Hot Sure Can Be Annoying!
Ugh, It’s Tough Being a Model
Ugh, I’m Too Skinny!
Ugh, People Keep Hitting on Me!
Achievements Ugh, Can You Believe They Included Me on This List?
Ugh, I Can’t Believe I Won an Award
Ugh, I’m So Successful
Ugh, I Can’t Believe I Was Mentioned in This Thing!
Performance at Work Ugh, How’d I Get Here??? How Is This My Life???
Money and Wealth Ugh, I Hate Having All This Money!
Intelligence Ugh, I’m a Genius
Personality Ugh, I’m So Humble!
Ugh, It’s Hard Being So Charitable!
Social Life Ugh, It’s So Weird Getting Recognized!
Ugh, I Hate People Wanting My Picture and Autograph All the Time
Ugh, I’m at an Exclusive Event!
Ugh, Being an Author Is Hard!
Miscellaneous Ugh, I Travel Too Much!
Table 6: Mapping Wittel’s humblebrag themes to Sezer et al.’s humblebrag categories.